# Systematic analysis of critical exponents in continuous dynamical phase transitions of weak noise theories

### v1

@@ I Introduction
Despite the significant advancements in nonequilibrium physics, we are still far from possessing a unified theory with a comparably comprehensive scope as equilibrium statistical mechanics. In equilibrium systems, statistical mechanics allows us to calculate thermodynamic variables by counting the number of microstates corresponding to a given macrostate, leading to well-defined quantities such as entropy and free energy. However, the powerful idea of state counting does not generally extend to nonequilibrium systems. In the 1970s, David Ruelle proposed shifting the focus from states to trajectories by introducing the idea of counting trajectories associated with time-integrated quantities, putting forward large deviation functions as nonequilibrium analogues of thermodynamic potentials [11, 12]. In what follows, we take advantage of this analogy.
@@
Some progress has been made in understanding DPTs in Langevin systems with weak noise. A weak noise Langevin system indicates that stochastic fluctuations are small compared to the deterministic forces, allowing the noise to be treated as a perturbation. This simplification makes it easier to obtain analytical and numerical results, as the dynamics can often be captured by focusing on rare but significant fluctuations that dominate the system’s behavior. For several important models, different approaches have successfully identified and validated the CEs characterizing these transitions [18, 19, 20, 21]. A prevalent method has been to emulate the Landau theory of phase transitions [22], which relies on symmetry considerations and the identification of an appropriate order parameter [23, 24, 25, 26]. However, the applicability of the Landau theory is limited if the relevant symmetries or the order parameter are difficult to identify. Interestingly, there are instances where different spontaneous symmetry breaking mechanisms have led to the same CEs [19], suggesting a greater degree of universality than previously expected. This has led to conjectures that the universality classes of DPTs might be broader and more interconnected than those of traditional thermodynamic phase transitions.
@@ II Setup and results
where $F[\phi]$ represents the deterministic force acting on the field, $G[\phi]$ characterizes the possibly field-dependent fluctuations, and $\xi(x,t)$ is standard Gaussian space-time white noise with $\left\langle\xi(x,t)\right\rangle=0$, $\left\langle\xi(x,t)\xi(x^{\prime},t^{\prime})\right\rangle=\delta(x-x^{\prime})\delta(t-t^{\prime})$. The functionals $F,G$ may include spatial derivative terms. The parameter $\varepsilon>0$ quantifies the relative strength of the stochastic noise compared to the deterministic dynamics. The path probability $P[\phi]$ of observing the path $\phi$ through space and time is given by
@@
In this work we consider $\varepsilon\ll 1$, where the discretization scheme becomes irrelevant to leading order in $\varepsilon$, cf. [30, 31, 32, 33] for detailed discussions of the impact of different stochastic integral conventions and path integral discretization schemes. Hence, crucially, the discretization choice is immaterial to both the form of the action $S[\phi]$ and to the CEs in the weak noise limit $\varepsilon\to 0$ studied here.
@@
Here, $\delta z=z_{\mathrm{c}}-z>0$, $I_{\text{s}}$ is the singular scaling function of the large deviation function, and $\lambda,\theta>0$ are the CEs. Note that the singular scaling function is expected to satisfy $\varepsilon^{\lambda}I_{\text{s}}(\delta z/\varepsilon^{\theta})\propto\varepsilon$ for $\delta z$ finite and $\varepsilon\downarrow 0$, which determines the scaling of $I_{\text{s}}(\mu\gg 1)$. The main goal of this work is providing a recipe to determine the CEs $\lambda,\theta$. It should be stressed that $\lambda,\theta$ are not the textbook CEs $\alpha,\beta,\gamma,\delta,\nu$ [13, 14, 15, 16, 17] since we are interested in the rate function $I$ itself, i.e. the analogue of the free energy density, instead of its order parameter associated derivatives. We postpone clarifying the relations between the CEs $\lambda,\theta$ and the textbook CEs to the summary in section VII.
Section IV provides a detailed recipe to extracting the CEs $\lambda,\theta$, the results of which are summarized in table 1 in the simplest case. It stands on two main legs, taking a page from the theory of second order thermodynamic phase transitions. The first leg assumes the Ginzburg criterion [14, 15], stating that the dominance of the fluctuations becomes comparable to the mean field analysis in the second derivative of the large deviation function in (8). The Ginzburg criterion implies the relation $\lambda=2\theta$ for a second-order DPT. The second leg recalls Le Chatelier’s principle [38, 39], or simply put, the stability analysis of the optimal fluctuation. As $\delta z\downarrow 0$, stability of the optimal fluctuation $\bar{\phi}$ as a local minimum is compromised by a set of subleading fluctuations. Accounting for the contribution of the destabilizing fluctuations allows to acquire the scaling form of $I_{\text{s}}$ when $\delta z\gg\varepsilon^{\theta}$, and in this limit the singular part of (8) becomes linear in $\varepsilon$. This step leads to a second relation between the CEs $\lambda,\theta$. Therefore, the two legs allow to find the CEs $\lambda,\theta$ explicitly. We name this procedure the Gaussian fluctuation method.
@@ C $L$ and $T$ infinite $2$ $1$
Before we tackle weak-noise Langevin systems, we start with a pedagogical example of the CW model at thermal equilibrium.
@@ III The Curie-Weiss model
$\simeq$ $g_{z}(m^{*})+\textrm{subleading terms},$
where $m^{*}=m^{*}(z)$ indicates the global minimizer of $g_{z}$. For $10<z<$ 1, i.e. in high temperatures, we have $m^{*}=0$, whereas at $z>1$, $m=0$ is a local maximum of $g_{z}$ instead, and $m^{*}\neq 0$ satisfies $m^{*}=\tanh\left(zm^{*}\right)$. This is a hallmark of a second order phase transition, with the critical point at $z_{\mathrm{c}}=1$ indicating a singularity in the free energy density $\mathfrak{f}_{N}(z)$.
@@ III.1 Finding the critical exponents
The singularity in the free energy is smoothed out when taking into account finite size corrections in $N$. To extract the CEs, one typically expands $g_{z}$ for $z\uparrow z_{\mathrm{c}}$ around $m^{*}=0$, leading to a Landau theory with $m$ being the order parameter. We avoid using here the ideas of the Landau theory altogether. Instead, we infer the CEs from the quadratic (Gaussian) expansion in $m$ only. Note that for the Landau theory, we would need to expand at least to order $m^{4}$, such that stability of the Landau theory is ensured.
@@
${\cal Z}_{N}(z)$ $=$ $\sqrt{\frac{Nz}{2\pi}}e^{N\log 2}\int_{-\infty}^{\infty}\mathrm{d}m\,e^{-\frac{1}{2}Nz(z_{\mathrm{c}}-z)m^{2}+O(m^{4})}$ (13)
$=$ $e^{N\log 2-\frac{1}{2}\log\left(z_{\mathrm{c}}-z\right)+O(1/N)}\,,$
taking into account only Gaussian fluctuations around the minimizer $m^{*}=0$. Clearly, the free energy in the large $N$ limit displays a logarithmic singularity as $z\uparrow z_{\mathrm{c}}$. One can write the free energy density
$\mathfrak{f}_{N}(z)=-\log 2+\frac{1}{2N}\log\delta z\,+O(1/N),$ (14)
where $\delta z=z_{\mathrm{c}}-z>0$. In the case of a continuous phase transition, the free energy density is composed of an analytic and a singular contribution:
@@
We can immediately infer from (14) that $\lambda=1$ here. To find $\theta$, we demand that the second derivative of the singular scaling function in (III.1) with respect to $z$ becomes comparable to the leading order contribution in the critical window, i.e. when we take $\delta zN^{\theta}=\text{const}$. This is just a restatement of the Ginzburg criterion, implying
@@
In the present example, this implies $\lambda=1,\theta=1/2$. To summarize, by evaluating just the Gaussian fluctuations around $m=0$ at fixed $z<z_{\text{c}}$ and using the Ginzburg criterion (16), we have found the CEs of the CW model. This is the content of the Gaussian fluctuation method.
@@ IV Classification of critical exponents
$\exp\left\{\varepsilon^{-1}I(z)\right\}$
@@
$\simeq\exp\bigg{\{}\varepsilon^{-1}I_{0}(z)\;-\frac{1}{2}\log\varepsilon$
@@
where $S_{2}[\delta\phi]=\int\mathrm{d}x\int\mathrm{d}t\,\delta\phi\,\left(\delta^{2}S_{\bar{\phi}}/\delta\phi^{2}\right)\delta\phi$ is the second variation of $S$ at $\bar{\phi}$. Notice that since $\bar{\phi}$ is the solution of a minimization problem, the first variation $S_{1}[\delta\phi]=\int\mathrm{d}x\int\mathrm{d}t\;\delta\phi\;\delta S_{\bar{\phi}}/\delta\phi$ vanishes at $\bar{\phi}$. The boundary conditions for the Gaussian fluctuations $\delta\phi$ in (IV) are such that
@@
where $\delta I(z)=I(z)-I_{0}(z)$. Importantly, the (possibly linearly transformed) Fourier decomposition diagonalizes the second variation, so that different modes can be treated independently. Here, $g_{z}$ is a non-negative function for $\delta z>0$, as we have assumed that $\bar{\phi}$ is indeed a true and nondegenerate local minimum.
Our goal is to rearrange the right hand side of (21) into the anomalous and subleading correction terms in (8). In order to do that, we stress that the Langevin equation (1) is typically a hydrodynamic equation. Hence, there are microscopic cutoffs in time and distance. This translates to having cutoffs for the Fourier frequencies $k_{\text{max}},\omega_{\text{max}}$. While it is impossible to infer the value of the cutoffs from the hydrodynamic theory, universality of the second order DPT states that the the singular scaling function $I_{\text{s}}$ cannot depend on these cutoffs. This is because the DPT involves a macroscopic change in the properties of the system due to fluctuations, which cannot be attributed to microscopic details. In simpler terms and for the CW toy example: the inter-spin coupling, i.e. the specific value of $J$ affected by the neighboring spins distance, cannot affect the CE values.
@@ IV.1 One continuous Fourier frequency component
Indeed, the contribution to the singular terms arises from the $k_{0}$ frequency. Moreover, the significant contribution of the $k_{0}$ frequency comes from the vicinity of $\omega$ around $\omega_{0}$, see figure 1. This suggests considering a Taylor expansion of $g_{z}(k_{0},\omega)$ around $\omega=\omega_{0}$, which is $g_{z}(k_{0},\omega)=A\left(\delta z\right)^{n}+B\left(\delta\omega\right)^{2}+C\left(\delta\omega\right)^{3}+\dots$. Here, the positivity of $g_{z}$ requires a vanishing linear term in $\delta\omega=\omega-\omega_{0}$, and $B>0$. Note that additional mixed terms $\propto(\delta z)^{n}\,\delta\omega$ and higher orders may appear in the Taylor expansion of $g_{z}(k_{0},\omega)$ for certain field theories, such as the KPZ example from section V, but these will not change the following argument, and hence we omit them here.
@@
$\delta I(z)=-\varepsilon\frac{2\pi}{\Delta\omega}\sqrt{\frac{A}{B}}\left(\delta z\right)^{n/2}+\varepsilon\mathcal{C}\left(k_{\text{max}},\omega_{\text{max}}\right).$ (26)
@@
To obtain the second equation, we use the second derivative test or the Ginzburg criterion as discussed above. Here, we demand that the second derivative of the contribution from $I_{\text{s}}$ with respect to $z$, will be comparable to the leading order contribution. This simply implies $\lambda=2\theta$. This description corresponds to a second order DPT. If we have a continuous DPT of order $G_{\mathrm{c}}$ instead, with $G_{\mathrm{c}}$ an integer larger then one, then we deduce $\lambda=G_{\mathrm{c}}\theta$. The two scaling rules lead to

### v3

@@ I Introduction
Despite the significant advancements in nonequilibrium physics, we are still far from possessing a unified theory with a comparably comprehensive scope as equilibrium statistical mechanics. In equilibrium systems, statistical mechanics allows us to calculate thermodynamic variables by counting the number of microstates corresponding to a given macrostate, leading to well-defined quantities such as entropy and free energy. However, the powerful idea of state counting does not generally extend to nonequilibrium systems. In the 1970s, David Ruelle proposed shifting the focus from states to trajectories by introducing the idea of counting trajectories associated with time-integrated quantities, putting forward large deviation functions as nonequilibrium analogs of thermodynamic potentials [11, 12]. In what follows, we take advantage of this analogy.
@@
Some progress has been made in understanding DPTs in Langevin systems with weak noise. A weak noise Langevin system indicates that stochastic fluctuations are small compared to the deterministic forces, allowing the noise to be treated as a perturbation. This simplification makes it easier to obtain analytical and numerical results, as the dynamics can often be captured by focusing on rare but significant fluctuations that dominate the system’s behavior. For several important models, different approaches have successfully identified and validated the CEs characterizing these transitions [18, 19, 20, 21]. A prevalent method has been to emulate the Landau theory of phase transitions [22], which relies on symmetry considerations and the identification of an appropriate order parameter [23, 24, 25, 26, 27]. However, the applicability of the Landau theory is limited if the relevant symmetries or the order parameter are difficult to identify. Interestingly, there are instances where different spontaneous symmetry breaking mechanisms have led to the same CEs [19], suggesting a greater degree of universality than previously expected. This has led to conjectures that the universality classes of DPTs might be broader and more interconnected than those of traditional thermodynamic phase transitions.
@@ II Setup and results
where $F[\phi]$ represents the deterministic force acting on the field, $G[\phi]$ characterizes the possibly field-dependent fluctuations, and $\xi(x,t)$ is standard Gaussian space-time white noise with $\left\langle\xi(x,t)\right\rangle=0$, $\left\langle\xi(x,t)\xi(x^{\prime},t^{\prime})\right\rangle=\delta(x-x^{\prime})\delta(t-t^{\prime})$. The functionals $F,G$ may include spatial derivative terms, as well as non-local terms [28]. The parameter $\varepsilon>0$ quantifies the relative strength of the stochastic noise compared to the deterministic dynamics. The path probability $P[\phi]$ of observing the path $\phi$ through space and time under the Langevin equation (1) in the Itô interpretation is given by
@@
In this work we consider $\varepsilon\ll 1$, where the discretization scheme becomes irrelevant to leading order in $\varepsilon$, cf. [32, 33, 34, 35] for detailed discussions of the impact of different stochastic integral conventions and path integral discretization schemes. Hence, crucially, the discretization choice is immaterial to both the form of the action $S[\phi]$ and to the CEs in the weak noise limit $\varepsilon\ll 1$ studied here.
@@
Here, $\delta z=z_{\mathrm{c}}-z>0$, $I_{\text{s}}$ is the singular scaling function of the large deviation function, and $\lambda,\theta>0$ are the CEs. Note that the singular scaling function is expected to satisfy $\varepsilon^{\lambda}I_{\text{s}}(\delta z/\varepsilon^{\theta})\propto\varepsilon$ for $\delta z$ finite and $\varepsilon\downarrow 0$, which determines the scaling of $I_{\text{s}}(\mu\gg 1)$. The main goal of this work is providing a recipe to determine the CEs $\lambda,\theta$. It should be stressed that $\lambda,\theta$ are not the textbook CEs $\alpha,\beta,\gamma,\delta,\nu$ [13, 14, 15, 16, 17] since we are interested in the rate function $I$ itself, i.e. the analog of the free energy density, instead of its order parameter associated derivatives. We postpone clarifying the relations between the CEs $\lambda,\theta$ and the textbook CEs to the summary in section VII.
Section IV provides a detailed recipe to extracting the CEs $\lambda,\theta$, the results of which are summarized in table 1 in the simplest case. It stands on two main legs, taking a page from the theory of second order thermodynamic phase transitions. The first leg assumes the Ginzburg criterion [14, 15], stating that  for a second-order DPT, the dominance of the fluctuations becomes comparable to the mean field analysis in the second derivative of the large deviation function in (8). The Ginzburg criterion implies the relation $\lambda=2\theta$ for a second-order DPT. The second leg recalls Le Chatelier’s principle [40, 41], or simply put, the stability analysis of the optimal fluctuation. As $\delta z\downarrow 0$, stability of the optimal fluctuation $\bar{\phi}$ as a local minimum is compromised by a set of subleading fluctuations. Accounting for the contribution of the destabilizing fluctuations allows to acquire the scaling form of $I_{\text{s}}$ when $\delta z\gg\varepsilon^{\theta}$, and in this limit the singular part of (8) becomes linear in $\varepsilon$. This step leads to a second relation between the CEs $\lambda,\theta$. Therefore, the two legs allow to find the CEs $\lambda,\theta$ explicitly. We name this procedure the Gaussian fluctuation method. We compare its range of applicability to the traditional Landau theory in table 2.
@@ C $L$ and $T$ infinite $2$ $1$
Before we tackle weak-noise Langevin systems with the described method, we start with a pedagogical example of the CW model at thermal equilibrium.
Gaussian Fluctuation Method Landau Theory
Conditions to use Knowledge of transition order $G_{\text{c}}\geq 2$, and assuming the unstable modes in the Gaussian terms vanish linearly in $\delta z$ ($n=1$) Identification of the order parameter, symmetries, and exact mean-field solution $\bar{\phi}(x,t)$
Predicts Critical exponents not directly related to the order parameter: $\theta,\lambda,\alpha,\nu$ All critical exponents: $\theta,\lambda,\alpha,\beta,\gamma,\delta,\nu$
Table 2: Comparison between Landau theory and the Gaussian fluctuation method presented here. The Landau theory provides a complete prediction of all critical exponents but requires an explicit mean-field solution, which is often challenging to obtain. The Gaussian fluctuation method, while limited in scope, is more broadly applicable, especially when the dynamical phase transition is complex and lacks an analytically tractable mean-field description.
@@ III The Curie-Weiss model
$\simeq$ $g_{z}(\bar{m})+\textrm{subleading terms},$
where $\bar{m}=\bar{m}(z)$ indicates the global minimizer of $g_{z}$. For $10<z<$ 1, i.e. in high temperatures, we have $\bar{m}=0$, whereas at $z>1$, $m=0$ is a local maximum of $g_{z}$ instead, and $\bar{m}\neq 0$ satisfies $\bar{m}=\tanh\left(z\bar{m}\right)$. This is a hallmark of a second order phase transition, with the critical point at $z_{\mathrm{c}}=1$ indicating a singularity in the free energy density $\mathfrak{f}_{N}(z)$.
@@ III.1 Finding the critical exponents
The singularity in the free energy is smoothed out when taking into account finite size corrections in $N$. To extract the CEs, one typically expands $g_{z}$ for $z\uparrow z_{\mathrm{c}}$ around $\bar{m}=0$, leading to a Landau theory with $m$ being the order parameter. We avoid using here the ideas of the Landau theory altogether. Instead, we infer the CEs from the quadratic (Gaussian) expansion in $m$ only. Note that for the Landau theory, we would need to expand at least to order $m^{4}$, such that stability of the Landau theory is ensured.
@@
${\cal Z}_{N}(z)$ $=$ $\sqrt{\frac{Nz}{2\pi}}e^{N\log 2}\int_{-\infty}^{\infty}\mathrm{d}m\,e^{-\frac{1}{2}Nz(z_{\mathrm{c}}-z)m^{2}+O(Nm^{4})}$ (13)
$=$ $e^{N\log 2-\frac{1}{2}\log\left(z_{\mathrm{c}}-z\right)}(1+O(1/N))\,,$
taking into account only Gaussian fluctuations around the minimizer $\bar{m}=0$. One can hence write the free energy density as
$\mathfrak{f}_{N}(z)=-\log 2+\frac{1}{2N}\log\delta z\,+O\left(1/N^{2}\right),$ (14)
where $\delta z=z_{\mathrm{c}}-z>0$, in this case. Clearly, the free energy in the large $N$ limit displays a logarithmic singularity as $z\uparrow z_{\mathrm{c}}$.
In the case of a continuous phase transition, the free energy density in the vicinity of the critical point is composed of an analytic and a singular contribution:
@@
Here, the singular scaling function $\mathfrak{f}_{\text{s}}$ can include contributions from the omitted higher-order terms in (14), which become comparable to the $O(1/N)$ term in the critical window. Nevertheless, even without knowing these terms explicitly, we still can immediately infer from (14) that $\lambda=1$ here. To find $\theta$, we demand that the second derivative of the singular scaling function in (III.1) with respect to $z$ becomes comparable to the leading order contribution in the critical window, i.e. when we take $\delta zN^{\theta}=\text{const}$. This is just a restatement of the Ginzburg criterion, implying
@@
In the present example, this implies $\lambda=1,\theta=1/2$. To summarize, by evaluating just the Gaussian fluctuations around $\bar{m}=0$ at fixed $z<z_{\text{c}}$ and using the Ginzburg criterion (16), we have found the CEs of the CW model. This is the content of the Gaussian fluctuation method.
@@ IV Classification of critical exponents
$\exp\left\{-\varepsilon^{-1}I(z)\right\}$
@@
$\simeq\exp\bigg{\{}-\varepsilon^{-1}I_{0}(z)\;-\frac{1}{2}\log\varepsilon$
@@
where $S_{2}[\delta\phi]=\int\mathrm{d}x\int\mathrm{d}t\,\delta\phi\,\left(\delta^{2}S_{\bar{\phi}}/\delta\phi^{2}\right)\delta\phi$ is the second variation of $S$ at $\bar{\phi}$. Notice that since $\bar{\phi}$ is the solution of a minimization problem, the first variation $S_{1}[\delta\phi]=\int\mathrm{d}x\int\mathrm{d}t\;\delta\phi\;\delta S_{\bar{\phi}}/\delta\phi$ vanishes at $\bar{\phi}$  [^1] For a non-local action, the explicit form of the expressions here would be different. Nevertheless, after applying the Fourier analysis in (21), the results presented in the following become consistent with a non-local action as well.. The boundary conditions for the Gaussian fluctuations $\delta\phi$ in (IV) are such that
@@
where $\delta I(z)=I(z)-I_{0}(z)$. Importantly, the (possibly linearly transformed) Fourier decomposition diagonalizes the second variation, so that different modes can be treated independently. Here, $g_{z}$ is a non-negative function for $\delta z>0$, as we have assumed that $\bar{\phi}$ is indeed a true and nondegenerate local minimum. We note that actually finding a diagonal decomposition of the second variation in practice can be difficult. While it is indeed possible in both examples we consider in sections V and VI, there are other cases where it may not be straightforward. We further comment on this point in section VII.
Our goal is to rearrange the right hand side of (21) into the anomalous and subleading correction terms in (8). In order to do that, we stress that the Langevin equation (1) is typically a hydrodynamic equation. Hence, there are microscopic cutoffs in time and distance. This translates to having cutoffs for the Fourier frequencies $k_{\text{max}},\omega_{\text{max}}$. While it is impossible to infer the value of the cutoffs from the hydrodynamic theory, universality of the second order DPT states that the singular scaling function $I_{\text{s}}$ cannot depend on these cutoffs. This is because the DPT involves a macroscopic change in the properties of the system due to fluctuations, which cannot be attributed to microscopic details. In simpler terms and for the CW toy example: the inter-spin coupling, i.e. the specific value of $J$ affected by the neighboring spins distance, cannot affect the CE values.
@@ IV.1 One continuous Fourier frequency component
Indeed, the contribution to the singular terms arises from the $k_{0}$ frequency. Moreover, the significant contribution of the $k_{0}$ frequency comes from the vicinity of $\omega$ around $\omega_{0}$, see Fig. 1. This suggests considering a Taylor expansion of $g_{z}(k_{0},\omega)$ around $\omega=\omega_{0}$, which is $g_{z}(k_{0},\omega)=A\left(\delta z\right)^{n}+B\left(\delta\omega\right)^{2}+C\left(\delta\omega\right)^{3}+\dots$. Here, the positivity of $g_{z}$ requires a vanishing linear term in $\delta\omega=\omega-\omega_{0}$, and $B>0$. Note that additional mixed terms $\propto(\delta z)^{n}\,\delta\omega$ and higher orders may appear in the Taylor expansion of $g_{z}(k_{0},\omega)$ for certain field theories, such as the KPZ example from section V, but these will not change the following argument, and hence we omit them here.
@@
$\delta I(z)=\varepsilon\frac{2\pi}{\Delta\omega}\sqrt{\frac{A}{B}}\left(\delta z\right)^{n/2}+\varepsilon\mathcal{C}\left(k_{\text{max}},\omega_{\text{max}}\right).$ (26)
@@
To obtain the second equation, we use the second derivative test or the Ginzburg criterion as discussed above. Here, we demand that the second derivative of the contribution from $I_{\text{s}}$ with respect to $z$, will be comparable to the leading order contribution. This simply implies $\lambda=2\theta$. This description corresponds to a second order DPT. If we have a continuous DPT of order $G_{\mathrm{c}}$ instead, with $G_{\mathrm{c}}$ an integer larger than one, then we deduce $\lambda=G_{\mathrm{c}}\theta$. The two scaling rules lead to